OpenStudy (_excusemyfrench._):

Oscar entered y= -0.4^x+3 into an online graphing tool. What function did he graph? A. y=(-0.4)^x+^3 B. y=(-0.4)^x+3 C. y=-(0.4)^x+^3 D. y=-(0.4)^x+3

OpenStudy (welshfella):

which one do you think it is?

OpenStudy (_excusemyfrench._):

maybe A..?

OpenStudy (tkhunny):

Maybe. Justify your response.

OpenStudy (welshfella):

does ^3 make sense to you?

OpenStudy (tkhunny):

Oscar entered no decimal. Why in the name of reason would it be C or D?

OpenStudy (mathmale):

Be aware that "Order of Operations" rules govern the evaluation of this expression. Are you familiar with these rules, or do you need a review?

OpenStudy (_excusemyfrench._):

sorry i typed that wrong. 0.4 is the decimal he entered

OpenStudy (mathmale):

Oh. Better double check everything before hitting the Enter key. That typo cost you, me and others precious time. Try again, from scratch. Look up order of operations rules, and then apply them to evaluate the expression.

OpenStudy (tkhunny):

The correct problem statement sure makes a difference. There may be multiple correct responses. The author of the question may not know that.

OpenStudy (welshfella):

do you think the negative 0.4 is taken to the power x or just the 0.4?

OpenStudy (_excusemyfrench._):

i think its taken to the x power

OpenStudy (solomonzelman):

then \(4^{\sqrt{2}}\) for instace will not have a value (it's not going to be odd or even, it won't be negative or positive)

OpenStudy (solomonzelman):

I meant\((-0.4)^{\sqrt{2}}\)

OpenStudy (solomonzelman):

Exponential function is continuous (because it is defined for all x), but for all irrational numbers, it will have gaps

OpenStudy (solomonzelman):

if you make the base negative

OpenStudy (mathmale):

Could we please focus on the simplest way or ways of looking at the problem in question, and save the rest for enrichment, later? Please?

OpenStudy (_excusemyfrench._):

i know the order of operations but i dont know how to apply it to the equation to get one of the answer choices

OpenStudy (tkhunny):

In some languages Unary Minus is higher precedence that Exponent. This may result in surprising errors or otherwise inconsistent results.

OpenStudy (_excusemyfrench._):

guys...im kinda confused..

OpenStudy (tkhunny):

And well you should be. This is a horribly written problem. What online graphing tool? Does it translate left-to-right or right-to-left r some other precedence rules? How does it treat Unary Minus? Does it even use 'x' and 'y'? Horrible, horrible problem statement. For the non-silly part, does "+^3" even mean anything?

OpenStudy (_excusemyfrench._):

^3 is an exponent..i dont know how to type an actual exponent.

OpenStudy (mathmale):

I'm going to assume that the following exactly represents the problem at hand: Oscar entered y= -0.4^x+3 into an online graphing tool. What function did he graph? A. y=(-0.4)^x+^3 B. y=(-0.4)^x+3 C. y=-(0.4)^x+^3 D. y=-(0.4)^x+3 Order of Ops rules require doing exponentiation first. Thus, 0.4^x is done first. The same rules require that mult. and div. be done next. Thus, 0.4^x becomes -0.4^x. All that remains to be done is to add 3 to -0.4^x. So, ExcuseMF, which answer seems to be the correct one?

OpenStudy (welshfella):

I believe if you entered it as you did in the question the sofware would take the 0.4 to the power x than convert it to negative

OpenStudy (mathmale):

If 3 is an exponent, it looks like we're back to square 1. ;( Could you please take a screenshot of the original problem and share that screen shot here?

OpenStudy (mathmale):

Welshie: "I believe if you entered it as you did in the question the sofware would take the 0.4 to the power x than convert it to negative" .... I agree fully.

OpenStudy (mathmale):

Keep in mind that the calculator is programmed to follow order of operations rules. So, whatever you type in is evaluated by the calculator according to those rules, unless you've made a glaring error in typing in the mathematical expression.

OpenStudy (mathmale):

I certainly don't want to make you feel bad, but again ask for a screen shot of the original problem. Having that in front of me would help me to help you write your problem statement correctly (including typing in exponents).

OpenStudy (tkhunny):

I just entered "y=0.4^x+3" in FooPlot and received "Invalid variable or function: y" I don't see that in the list of choices. Hideous question.

OpenStudy (tkhunny):

DewsMos gives \(y = -(0.4)^{x} + 3\)

OpenStudy (tkhunny):

Mathway gives \(y = -(0.4)^{x+3}\)

OpenStudy (tkhunny):

I don't want to make YOU feel bad, either, but I do want to make the author of the question feel bad.

OpenStudy (welshfella):

yes it is a bad question

OpenStudy (mathmale):

Not much chance of that happening, tk, but otherwise I'm with you.

OpenStudy (mathmale):

However....... we all really do need to know exactly how a calculator will respond to what we type into it. So I think this question is still worth the thought.

OpenStudy (_excusemyfrench._):

okay. not trying to make this anymore complicated than it already is..but i dont know how to screenshot..and my teacher has the question's equation exactly the same as i typed it. the only difference is the A and C have the number 3 as an actual exponent.:/

OpenStudy (_excusemyfrench._):

tk does that mean that theres no answer to the question..?

OpenStudy (mathmale):

Thank goodness that something are standardized (as are most calculator responses to input). Frenchie: could you possibly take a cell phone photo of this problem, upload it to your computer, and from your computer upload it to OpenStudy? It'd be worth your while to ask whether your computer is capable of taking screen shots and if it is, to learn how to do them. Perhaps you and I could work on that later.

OpenStudy (mathmale):

Unfortunately, I need to get off my computer for the present. We've spent a lot of time on this problem, perhaps at the expense of others that you might understand more readily. Might be time to move on now, and return to this question only if you find you have the time and energy to do so. Excuse me, I'm saying good-bye now.

OpenStudy (_excusemyfrench._):

well...okay im gonna close it. thanks for trying guys..srry..wasted your time

OpenStudy (mathmale):

We all learned something. Just hope you'll take me up on my offer to help you capture screen shots and share them.

OpenStudy (_excusemyfrench._):

okay one sec

OpenStudy (welshfella):

I entered the equation on Desmos graphing calculator and my own graphical computer and got the same curve Now i'm trying option D which I think is the correct one.

OpenStudy (welshfella):

Yes thats it.

OpenStudy (welshfella):

The calculators follow the PEMDAS rules of operations

OpenStudy (welshfella):

it takes 3 as not part of the exponent To make it part of the exponent you would have to enter (x + 3) after the '^'

OpenStudy (welshfella):

do you follow that?

OpenStudy (_excusemyfrench._):

yeah i think so

OpenStudy (welshfella):

the software takes just x to be the exponent unless you put the x+3 in parentheses

OpenStudy (_excusemyfrench._):

can i use a scientific calculator?

OpenStudy (welshfella):

you need a graphing calculator

OpenStudy (_excusemyfrench._):

oh okay

OpenStudy (welshfella):

what you could do is replace the x by say 2 and enter the formula in the question and the formula in option D . they both should give the same answer.

OpenStudy (welshfella):

-0.4^2 + 3 and -(0.4)^2 + 3

OpenStudy (_excusemyfrench._):

okay so put -0.4^2+3 in desmos & -(0.4)^2+3?

OpenStudy (welshfella):

right

OpenStudy (_excusemyfrench._):

okay i did that

OpenStudy (welshfella):

they should both give you the same answer -oh does desmos do calculations as well??

OpenStudy (welshfella):

Yes it does Did they both give the same result?

OpenStudy (_excusemyfrench._):

2.84?

OpenStudy (welshfella):

yes

OpenStudy (_excusemyfrench._):

Yep they have the same answer

OpenStudy (welshfella):

yep a scientific calculator will give the same answer too ( well should do!) so D is definitely the correct answer.

OpenStudy (_excusemyfrench._):

omg thank you soo much! you made it so easy. fan & medal

OpenStudy (welshfella):

ty yw

OpenStudy (tkhunny):

It's an okay question as far as Order of Operations goes, but not as it is phrased with the online business.